OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2). (4 points) Question 5 options: 1) y = -5x + 13 2) y = -1/5 x + 7/5 3) y = 1/5 x - 7/5 4) y = 5x - 17 *Help please

OpenStudy (anonymous):

so if it is perpendicular, your slope should be -1/5 so 2.

OpenStudy (anonymous):

Thank you @melodysim

jimthompson5910 (jim_thompson5910):

The slope of y = 5x + 4 is 5 or 5/1 Flip the fraction 5/1 ---> 1/5 Flip the sign: +1/5 ---> -1/5 So the perpendicular slope is -1/5 and only choice 2) has this slope

OpenStudy (anonymous):

Oh okay gotcha ! The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, -2). (4 points) Question 4 options: 1) y = -2x + 4 2) y = x - 3) y = x - 4) y = 2x - 8